Can someone help me please? I'm having a real hard time with this.
A student makes a $9.50 purchase at the bookstore with a 20 dollar bill. The store has no bills and gives the change in quarters and 50-cent pieces. There are 30 coins in all.How many of each kind are there? Thanks alot.

Answer by stanbon(75887)

A student makes a $9.50 purchase at the bookstore with a 20 dollar bill. The store has no bills and gives the change in quarters and 50-cent pieces. There are 30 coins in all.How many of each kind are there?
----------------
Change = $10.50
------------------
Equations:
Quantity Eq.::: q + f = 30
Value Eq;;;;;;25q + 50f = 1050 cents
---------------------------------------------
Solve by substitution:
q = 30-f
Substitute to get: 25(30-f) + 50f = 1050
750 - 25f+ 50f = 1050
25f = 300
f = 12 (# of 50-cent pieces)
-----------------------
Substitute into q + f = 30 to solve for "q":
q + 12 = 30
q = 18 (# of quarters in the change)
================================================
Cheers,
Stan H.